OpenStudy (anonymous):

A collection of quarters and nickels is worth $13.70. There are 106 coins in all. How many of each coin are there? quarters nickels

OpenStudy (anonymous):

if I fail this test my mom will hurt me

OpenStudy (anonymous):

8 and 19

OpenStudy (anonymous):

let the shorter piece = x longer piece = 2x + 3 2x+ 3 + x= 27 3x = 24 you can find x and then longer piece

OpenStudy (anonymous):

so the long piecis is 19

OpenStudy (anonymous):

noooo

OpenStudy (anonymous):

Of course it is

OpenStudy (anonymous):

Admission for a movie is $8 for children and $12 for adults. On a certain day, 3200 people attend the movie theater and $33,040 is collected. How many children and how many adults attended the movies?

OpenStudy (anonymous):

can anyone help I have to pass this because I do k12 and I only take it once im begging please

OpenStudy (anonymous):

let say there are x children and y adults 1. x+y = 3200 2. 8x +12y = 33040 3. x= 3200 - y put value of x from 3. to 2, equation

OpenStudy (anonymous):

wait so what

OpenStudy (anonymous):

1340 1860 is that it

OpenStudy (anonymous):

Yups

OpenStudy (anonymous):

A collection of quarters and nickels is worth $13.70. There are 106 coins in all. How many of each coin are there?

OpenStudy (anonymous):

can anyone help

OpenStudy (anonymous):

Let quarte = x and nickels = y 1. x+ y = 106 2. 25x + 5y = 1370 3. x = 106 - y Put the value of x from 3. to 2.

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

42 64 i dnt understand

OpenStudy (anonymous):

what part?

OpenStudy (anonymous):

is this correct

OpenStudy (anonymous):

Put these values you got in 2. equation and check whether it is correct

OpenStudy (anonymous):

the test has 5 questions i cant submit it yet

OpenStudy (anonymous):

oh my gd sangya saved my life yay

OpenStudy (anonymous):

You solved it. Congrats

OpenStudy (anonymous):

Airplane A takes off and ascends at a rate of 3000 ft/min. At the same time, Airplane B, which is flying at 30,000 ft, begins to descend at a rate of 2500 ft/min. Let h represent the height off the ground in feet, and let t represent time in minutes. Which system models this situation?